# Constant-Size Cryptographic Evidence Structures for Regulated AI Workflows

### v1

@@ Abstract
This paper introduces constant-size cryptographic evidence structures, a general abstraction for representing verifiable audit evidence for AI workflows in regulated environments. Each evidence item is a fixed-size tuple of cryptographic fields, designed to (i) provide strong binding to workflow events and configurations, (ii) support constant-size storage and uniform verification cost per event, and (iii) compose cleanly with hash-chain and Merkle-based audit constructions. We formalize a simple model of regulated AI workflows, define syntax and algorithms for evidence structures, and articulate security goals such as audit integrity and non-equivocation.
We present a generic hash-and-sign construction that instantiates this abstraction using a collision-resistant hash function and a standard digital signature scheme. We then show how to integrate the construction with hash-chained logs, Merkle-tree anchoring, and (optionally) ted execution environments, and we analyze the asymptotic complexity of evidence generation and verification. Finally, we implement a prototype library and report microbenchmark results on commodity hardware, demonstrating that the per-event overhead of constant-size evidence is small and predictable. The design is informed by industrial experience with regulated AI systems at Codebat Technologies Inc., while the paper focuses on the abstraction, algorithms, and their security and performance characteristics. This work aims to provide a foundation for standardized audit mechanisms in regulated AI, with implications for clinical trial management, pharmaceutical compliance, and medical AI governance.
@@ 1 Introduction
Secure logging and audit-trail systems have been studied extensively [schneier1999secure, crosby2009efficient, bellare1997forward], as have hash-chain and Merkle-tree-based data structures [merkle1980protocols, haber1991time]. More recently, blockchain-based transparency logs [laurie2013certificate, nakamoto2008bitcoin] and ted execution environments (TEEs) [costan2016intel, sabt2015ted] have been used to build tamper-evident records for security-critical operations. However, these approaches typically operate on variable-length or application-specific records and do not treat a fixed-size, hardware-friendly evidence representation as a first-class abstraction.
@@
    We present a generic hash-and-sign construction that instantiates the abstraction using a collision-resistant hash function and a digital signature scheme, and we give informal security arguments for audit integrity and non-equivocation.
@@
    We discuss how the construction composes with hash-chained logs, Merkle-tree anchoring, and optional TEE-based environment binding, and we analyze the asymptotic complexity of evidence generation and verification.
@@ 2.3 Design Objectives
Evidence items should integrate cleanly with existing audit structures such as hash chains and Merkle trees [merkle1980protocols], so that large collections of items can be anchored succinctly.
@@ 4.1 Components
Let $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$ be a collision-resistant hash function, and let $(\mathsf{KeyGen}^{\mathsf{sig}},\mathsf{Sign},\mathsf{Verify}^{\mathsf{sig}})$ be a digital signature scheme that is existentially unforgeable under chosen-message attacks (EUF-CMA). We assume that the system designer provides a family of deterministic encoding functions
@@
for each index $i\in I$, where $\mathcal{E}$ is the set of events.
@@ 4.2 Algorithms in Pseudocode
4.3 Informal Security Discussion
Assuming $H$ is collision-resistant and the signature scheme is EUF-CMA secure:
Binding to events.
If an adversary can produce two different events $E\neq E^{\prime}$ and a single evidence item $\mathsf{ev}$ such that both pairs $(E,\mathsf{ev})$ and $(E^{\prime},\mathsf{ev})$ verify under the same key, then either:
  • there exists an index $i$ with $H(\phi_{i}(E))=H(\phi_{i}(E^{\prime}))$ and $\phi_{i}(E)\neq\phi_{i}(E^{\prime})$, giving a collision in $H$, or
  • the adversary can forge a signature on $\mathsf{ev}$ without querying $\mathsf{Sign}$ on that message.
Both events contradict our assumptions.
Tamper detection.
If an adversary modifies any field of a stored evidence item $\mathsf{ev}$, then either the recomputed hashes will not match the encodings of the underlying event, or the original signature will fail to verify. Any successful undetected modification therefore implies either a hash collision or a signature forgery.
Chain integrity and non-equivocation.
When evidence items are linked with Algorithm 3 or with a Merkle tree, an adversary who attempts to remove, re-order, or equivocate about items must either:
  • find collisions in the hash function to produce two distinct sequences with the same chain tip or root, or
  • present two sets of items with different tips/roots and claim they correspond to the same anchored value.
In the first case, this contradicts collision resistance; in the second, the inconsistency is observable to an external auditor that tracks anchored values [haber1991time, laurie2013certificate].

### v2

@@ Abstract
This paper introduces constant-size cryptographic evidence structures, a general abstraction for representing verifiable audit evidence for AI workflows in regulated environments. Each evidence item is a fixed-size tuple of cryptographic fields, designed to (i) provide strong binding to workflow events and configurations, (ii) support constant-size storage and uniform verification cost per event, and (iii) compose cleanly with hash-chain and Merkle-based audit constructions. We formalize a simple model of regulated AI workflows, define syntax and algorithms for evidence structures, and formalize security properties—evidence binding, tamper detection, and non-equivocation—via game-based definitions, proving that our construction satisfies them under standard assumptions (collision-resistant hashing and EUF-CMA signatures).
We present a generic hash-and-sign construction that instantiates this abstraction using a collision-resistant hash function and a standard digital signature scheme. We then show how to integrate the construction with hash-chained logs, Merkle-tree anchoring, and (optionally) trusted execution environments, and we analyze the asymptotic complexity of evidence generation and verification. Finally, we implement a prototype library and report microbenchmark results on commodity hardware, demonstrating that the per-event overhead of constant-size evidence is small and predictable. The design is informed by industrial experience with regulated AI systems at Codebat Technologies Inc., while the paper focuses on the abstraction, algorithms, and their security and performance characteristics. This work aims to provide a foundation for standardized audit mechanisms in regulated AI, with implications for clinical trial management, pharmaceutical compliance, and medical AI governance.
Keywords: cryptographic evidence structures, evidence binding, game-based security proofs, regulated AI workflows, hash-and-sign construction, tamper detection
@@ 1 Introduction
Secure logging and audit-trail systems have been studied extensively [22, 7, 4], as have hash-chain and Merkle-tree-based data structures [15, 10]. More recently, blockchain-based transparency logs [13, 17] and trusted execution environments (TEEs) [6, 21] have been used to build tamper-evident records for security-critical operations. However, these approaches typically operate on variable-length or application-specific records and do not treat a fixed-size, hardware-friendly evidence representation as a first-class abstraction.
@@
    We present a generic hash-and-sign construction that instantiates the abstraction using a collision-resistant hash function and a digital signature scheme.
@@
    We formalize three security properties (evidence binding, tamper detection, non-equivocation) via game-based definitions and prove that each holds under standard assumptions with tight reductions.
@@
    We discuss how the construction composes with hash-chained logs, Merkle-tree anchoring, and optional TEE-based environment binding, and analyze the asymptotic complexity of evidence generation and verification.
  6. 6.
@@ 2.3 Design Objectives
Evidence items should integrate cleanly with existing audit structures such as hash chains and Merkle trees [15], so that large collections of items can be anchored succinctly.
@@ 4.1 Components
Let $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$ be a collision-resistant hash function, and let $(\mathsf{KeyGen}^{\mathsf{sig}},\mathsf{Sign},\mathsf{Verify}^{\mathsf{sig}})$ be a digital signature scheme that is existentially unforgeable under chosen-message attacks (EUF-CMA) [9, 18]. We assume that the system designer provides a family of deterministic encoding functions
@@
for each index $i\in I$, where $\mathcal{E}$ is the set of events. We require that the encoding family $\{\phi_{i}\}_{i\in I}$ is collectively injective: for any two distinct events $E\neq E^{\prime}$, there exists at least one index $i\in I$ such that $\phi_{i}(E)\neq\phi_{i}(E^{\prime})$. This ensures that distinct events produce differing inputs to the hash function in at least one field, which is essential for the binding property established in Section 5.
@@ 4.2 Algorithms in Pseudocode
5 Security Analysis
We now formalize the security guarantees of the hash-and-sign evidence construction. We define three security properties via game-based experiments [3, 5] and prove that the construction satisfies each under standard cryptographic assumptions [12]: collision resistance of the hash function [20] and existential unforgeability of the signature scheme [9]. Throughout this section, $\lambda$ denotes the security parameter, $H\colon\{0,1\}^{*}\to\{0,1\}^{\lambda}$ denotes a collision-resistant hash function, and $(\mathsf{KeyGen}^{\mathsf{sig}},\mathsf{Sign},\mathsf{Verify}^{\mathsf{sig}})$ denotes an EUF-CMA-secure signature scheme.
5.1 Security Definitions
We define three game-based experiments that capture the core security goals of constant-size evidence structures.
Definition 5.1 (Evidence Binding).
Let $\Pi$ be an evidence structure with field count $k$. The evidence-binding experiment $\mathbf{Exp}^{\mathrm{Bind}}_{\Pi,\mathcal{A}}(\lambda)$ proceeds as follows:
  1. 1.
    Run $\mathsf{pp}\leftarrow\mathsf{Setup}(1^{\lambda})$ and $(sk,pk)\leftarrow\mathsf{KeyGen}(1^{\lambda})$.
  2. 2.
    The adversary $\mathcal{A}$ receives $\mathsf{pp}$ and $pk$, and has oracle access to $\mathsf{Generate}(\mathsf{pp},sk,\cdot)$.
  3. 3.
    $\mathcal{A}$ outputs $(E,E^{\prime},\mathsf{ev},\sigma)$.
  4. 4.
    The experiment outputs $1$ if and only if $E\neq E^{\prime}$, $\mathsf{Verify}(\mathsf{pp},pk,E,\mathsf{ev},\sigma)=\mathsf{accept}$, and $\mathsf{Verify}(\mathsf{pp},pk,E^{\prime},\mathsf{ev},\sigma)=\mathsf{accept}$.
The evidence-binding advantage of $\mathcal{A}$ is $\mathbf{Adv}^{\mathrm{Bind}}_{\Pi,\mathcal{A}}(\lambda)=\Pr[\mathbf{Exp}^{\mathrm{Bind}}_{\Pi,\mathcal{A}}(\lambda)=1]$.
Definition 5.2 (Tamper Detection — Unforgeability).
Let $\Pi$ be an evidence structure. The tamper-detection experiment $\mathbf{Exp}^{\mathrm{TD}}_{\Pi,\mathcal{A}}(\lambda)$ proceeds as follows:
  1. 1.
    Run $\mathsf{pp}\leftarrow\mathsf{Setup}(1^{\lambda})$ and $(sk,pk)\leftarrow\mathsf{KeyGen}(1^{\lambda})$.
  2. 2.
    The adversary $\mathcal{A}$ receives $\mathsf{pp}$ and $pk$, and has oracle access to $\mathsf{Generate}(\mathsf{pp},sk,\cdot)$. Let $\mathcal{Q}$ be the set of events queried to $\mathsf{Generate}$.
  3. 3.
    $\mathcal{A}$ outputs $(E^{*},\mathsf{ev}^{*},\sigma^{*})$.
  4. 4.
    The experiment outputs $1$ if and only if $E^{*}\notin\mathcal{Q}$ and $\mathsf{Verify}(\mathsf{pp},pk,E^{*},\mathsf{ev}^{*},\sigma^{*})=\mathsf{accept}$.
The tamper-detection advantage of $\mathcal{A}$ is $\mathbf{Adv}^{\mathrm{TD}}_{\Pi,\mathcal{A}}(\lambda)=\Pr[\mathbf{Exp}^{\mathrm{TD}}_{\Pi,\mathcal{A}}(\lambda)=1]$.
Definition 5.3 (Non-Equivocation — Chain Integrity).
Let $\Pi$ be an evidence structure with hash-chain linking (Algorithm 3). The non-equivocation experiment $\mathbf{Exp}^{\mathrm{NE}}_{\Pi,\mathcal{A}}(\lambda)$ proceeds as follows:
  1. 1.
    Run $\mathsf{pp}\leftarrow\mathsf{Setup}(1^{\lambda})$.
  2. 2.
    The adversary $\mathcal{A}$ receives $\mathsf{pp}$.
  3. 3.
    $\mathcal{A}$ outputs two sequences of evidence items $S=(\mathsf{ev}_{1},\dots,\mathsf{ev}_{n})$ and $S^{\prime}=(\mathsf{ev}^{\prime}_{1},\dots,\mathsf{ev}^{\prime}_{m})$.
  4. 4.
    The experiment outputs $1$ if and only if $S\neq S^{\prime}$ and $\mathsf{Link}(\mathsf{pp},S)=\mathsf{Link}(\mathsf{pp},S^{\prime})$.
The non-equivocation advantage of $\mathcal{A}$ is $\mathbf{Adv}^{\mathrm{NE}}_{\Pi,\mathcal{A}}(\lambda)=\Pr[\mathbf{Exp}^{\mathrm{NE}}_{\Pi,\mathcal{A}}(\lambda)=1]$.
Remark.
Evidence binding (Definition 5.1) is analogous to the binding property of cryptographic commitment schemes: a single evidence item cannot serve as valid evidence for two distinct events. Together with tamper detection (Definition 5.2), these two properties jointly ensure what we informally call audit integrity: binding guarantees uniqueness of the event-to-evidence mapping, while tamper detection guarantees that valid evidence can only originate from honest executions.
5.2 Security Theorems
We now prove that the hash-and-sign construction from Section 4 satisfies all three security properties under standard assumptions.
Theorem 5.4 (Evidence Binding).
For any PPT adversary $\mathcal{A}$ against the evidence-binding experiment of the hash-and-sign evidence construction with field count $k$, there exists a PPT adversary $\mathcal{B}$ against the collision resistance of $H$ such that
$\mathbf{Adv}^{\mathrm{Bind}}_{\Pi,\mathcal{A}}(\lambda)\;\leq\;k\cdot\mathbf{Adv}^{\mathrm{CR}}_{H,\mathcal{B}}(\lambda).$
Proof.
Suppose $\mathcal{A}$ succeeds in the evidence-binding experiment with non-negligible probability. Then $\mathcal{A}$ outputs $(E,E^{\prime},\mathsf{ev},\sigma)$ with $E\neq E^{\prime}$ such that $\mathsf{Verify}(\mathsf{pp},pk,E,\mathsf{ev},\sigma)=\mathsf{accept}$ and $\mathsf{Verify}(\mathsf{pp},pk,E^{\prime},\mathsf{ev},\sigma)=\mathsf{accept}$.
Parse $\mathsf{ev}=(f_{0},\dots,f_{k-1})$. Because both verification checks pass with the same evidence item $\mathsf{ev}$, we have $f_{i}=H(\phi_{i}(E))$ and $f_{i}=H(\phi_{i}(E^{\prime}))$ for every $i\in\{0,\dots,k-1\}$. That is,
$H(\phi_{i}(E))=H(\phi_{i}(E^{\prime}))\quad\text{for all }i\in\{0,\dots,k-1\}.$
Since $E\neq E^{\prime}$, the encoding functions $\{\phi_{i}\}$ are collectively injective on distinct events (Section 4), so there exists at least one index $i^{*}$ with $\phi_{i^{*}}(E)\neq\phi_{i^{*}}(E^{\prime})$. For this index, the pair $(\phi_{i^{*}}(E),\;\phi_{i^{*}}(E^{\prime}))$ is a collision in $H$.
We construct a collision-finding adversary $\mathcal{B}$ as follows. $\mathcal{B}$ receives a hash function $H$ (sampled from the family), picks a random index $i\xleftarrow{\mathdollar}\{0,\dots,k-1\}$, and runs $\mathcal{A}$ faithfully, simulating the experiment using $H$. When $\mathcal{A}$ outputs $(E,E^{\prime},\mathsf{ev},\sigma)$, adversary $\mathcal{B}$ outputs $(\phi_{i}(E),\;\phi_{i}(E^{\prime}))$.
If $\mathcal{A}$ succeeds, then a collision exists at some index $i^{*}$, and $\mathcal{B}$’s random guess $i=i^{*}$ succeeds with probability at least $1/k$. Therefore,
$\mathbf{Adv}^{\mathrm{CR}}_{H,\mathcal{B}}(\lambda)\;\geq\;\frac{1}{k}\cdot\mathbf{Adv}^{\mathrm{Bind}}_{\Pi,\mathcal{A}}(\lambda),$
which gives the claimed bound. ∎
Theorem 5.5 (Tamper Detection).
For any PPT adversary $\mathcal{A}$ against the tamper-detection experiment of the hash-and-sign evidence construction with field count $k$, there exist PPT adversaries $\mathcal{B}_{1}$ against the collision resistance of $H$ and $\mathcal{B}_{2}$ against the EUF-CMA security of the signature scheme such that
$\mathbf{Adv}^{\mathrm{TD}}_{\Pi,\mathcal{A}}(\lambda)\;\leq\;k\cdot\mathbf{Adv}^{\mathrm{CR}}_{H,\mathcal{B}_{1}}(\lambda)\;+\;\mathbf{Adv}^{\mathrm{EUF\text{-}CMA}}_{\mathsf{Sig},\mathcal{B}_{2}}(\lambda).$
Proof.
Suppose $\mathcal{A}$ outputs $(E^{*},\mathsf{ev}^{*},\sigma^{*})$ with $E^{*}\notin\mathcal{Q}$ and $\mathsf{Verify}(\mathsf{pp},pk,E^{*},\mathsf{ev}^{*},\sigma^{*})=\mathsf{accept}$. We consider two cases.
Case 1: There exists a query $E\in\mathcal{Q}$ whose $\mathsf{Generate}$ output $(\mathsf{ev},\sigma)$ satisfies $\mathsf{ev}=\mathsf{ev}^{*}$.
Since $E^{*}\neq E$ (because $E^{*}\notin\mathcal{Q}$) and both $(E,\mathsf{ev})$ and $(E^{*},\mathsf{ev}^{*})$ pass verification with the same evidence item, the argument from Theorem 5.4 applies: there exists an index $i^{*}$ at which $\phi_{i^{*}}(E)\neq\phi_{i^{*}}(E^{*})$ yet $H(\phi_{i^{*}}(E))=H(\phi_{i^{*}}(E^{*}))$, yielding a collision in $H$. Adversary $\mathcal{B}_{1}$ guesses $i^{*}$ at random and succeeds with probability at least $1/k$.
Case 2: No query $E\in\mathcal{Q}$ produced an evidence item equal to $\mathsf{ev}^{*}$.
In this case, $\mathsf{ev}^{*}$ was never signed by the $\mathsf{Generate}$ oracle, yet $\sigma^{*}$ is a valid signature on $\mathsf{ev}^{*}$. We construct an EUF-CMA adversary $\mathcal{B}_{2}$ as follows: $\mathcal{B}_{2}$ receives a verification key $pk$ and access to a signing oracle $\mathsf{Sign}_{sk}(\cdot)$. It simulates $\mathsf{Setup}$ honestly (choosing $H$ and the encoding functions), and answers $\mathcal{A}$’s $\mathsf{Generate}$ queries by computing $\mathsf{ev}$ from $H$ and the encoding functions, then calling the signing oracle on $\mathsf{ev}$. When $\mathcal{A}$ outputs $(E^{*},\mathsf{ev}^{*},\sigma^{*})$, adversary $\mathcal{B}_{2}$ outputs $(\mathsf{ev}^{*},\sigma^{*})$ as its forgery.
Since no oracle query produced $\mathsf{ev}^{*}$, this is a valid EUF-CMA forgery. Combining both cases yields the stated bound. ∎
Theorem 5.6 (Non-Equivocation).
For any PPT adversary $\mathcal{A}$ against the non-equivocation experiment of the hash-chain linking construction (Algorithm 3), there exists a PPT adversary $\mathcal{B}$ against the collision resistance of $H$ such that
$\mathbf{Adv}^{\mathrm{NE}}_{\Pi,\mathcal{A}}(\lambda)\;\leq\;n\cdot\mathbf{Adv}^{\mathrm{CR}}_{H,\mathcal{B}}(\lambda),$
where $n=\max(|S|,|S^{\prime}|)$ is the maximum sequence length output by $\mathcal{A}$.
Proof.
Suppose $\mathcal{A}$ outputs two distinct sequences $S=(\mathsf{ev}_{1},\dots,\mathsf{ev}_{n})$ and $S^{\prime}=(\mathsf{ev}^{\prime}_{1},\dots,\mathsf{ev}^{\prime}_{m})$ such that $\mathsf{Link}(\mathsf{pp},S)=\mathsf{Link}(\mathsf{pp},S^{\prime})$, i.e., the chain tips are equal: $\ell_{n}=\ell^{\prime}_{m}$, where
$\ell_{0}=\ell^{\prime}_{0}=0^{\lambda},\quad\ell_{j}=H(\ell_{j-1}\parallel\mathsf{ev}_{j}),\quad\ell^{\prime}_{j}=H(\ell^{\prime}_{j-1}\parallel\mathsf{ev}^{\prime}_{j}).$
Since $S\neq S^{\prime}$, we consider two sub-cases.
Sub-case (a): $n=m$. The sequences have the same length but differ in at least one position. Walk backward from the tip: at $j=n$ we have $\ell_{n}=\ell^{\prime}_{n}$. Let $j^{*}$ be the largest index such that $\mathsf{ev}_{j^{*}}\neq\mathsf{ev}^{\prime}_{j^{*}}$ or $\ell_{j^{*}-1}\neq\ell^{\prime}_{j^{*}-1}$ (or both). Then
$H(\ell_{j^{*}-1}\parallel\mathsf{ev}_{j^{*}})=\ell_{j^{*}}=\ell^{\prime}_{j^{*}}=H(\ell^{\prime}_{j^{*}-1}\parallel\mathsf{ev}^{\prime}_{j^{*}}),$
and at least one of $\ell_{j^{*}-1}\neq\ell^{\prime}_{j^{*}-1}$ or $\mathsf{ev}_{j^{*}}\neq\mathsf{ev}^{\prime}_{j^{*}}$ holds, so $(\ell_{j^{*}-1}\parallel\mathsf{ev}_{j^{*}})\neq(\ell^{\prime}_{j^{*}-1}\parallel\mathsf{ev}^{\prime}_{j^{*}})$. This is a collision in $H$.
Sub-case (b): $n\neq m$. Without loss of generality, $n>m$. Define the augmented sequences $\hat{\ell}_{j}=\ell_{j}$ for $j=0,\dots,n$ and $\hat{\ell}^{\prime}_{j}=\ell^{\prime}_{j}$ for $j=0,\dots,m$ with $\hat{\ell}^{\prime}_{j}=\ell^{\prime}_{m}$ for $j=m+1,\dots,n$ (i.e., the shorter chain is padded with its final value). We have $\hat{\ell}_{0}=\hat{\ell}^{\prime}_{0}=0^{\lambda}$ and $\hat{\ell}_{n}=\ell_{n}=\ell^{\prime}_{m}=\hat{\ell}^{\prime}_{n}$. Consider the sequence of equalities $\hat{\ell}_{j}\stackrel{{\scriptstyle?}}{{=}}\hat{\ell}^{\prime}_{j}$ for $j=0,\dots,n$: they agree at $j=0$ and $j=n$. If they agree at every $j$, then in particular $\ell_{m}=\hat{\ell}_{m}=\hat{\ell}^{\prime}_{m}=\ell^{\prime}_{m}$ and $\ell_{m+1}=H(\ell_{m}\parallel\mathsf{ev}_{m+1})$ whereas $\hat{\ell}^{\prime}_{m+1}=\ell^{\prime}_{m}=\ell_{m}$. Since $n>m$ and the hash output is $\lambda$ bits, this forces $H(\ell_{m}\parallel\mathsf{ev}_{m+1})=\ell_{m}$, a collision in $H$ (unless $\ell_{m}\parallel\mathsf{ev}_{m+1}=\ell_{m}$, which is impossible since concatenation strictly extends the input). If they disagree at some index, let $j^{*}$ be the largest index where $\hat{\ell}_{j^{*}}\neq\hat{\ell}^{\prime}_{j^{*}}$ but $\hat{\ell}_{j^{*}+1}=\hat{\ell}^{\prime}_{j^{*}+1}$; then the hash inputs at step $j^{*}+1$ differ but outputs coincide, yielding a collision in $H$.
In both sub-cases, a collision in $H$ exists at some step $j^{*}\in\{1,\dots,\max(n,m)\}$. We construct $\mathcal{B}$ as follows: $\mathcal{B}$ picks $j\xleftarrow{\mathdollar}\{1,\dots,\max(n,m)\}$, runs $\mathcal{A}$, computes both chains, and outputs the hash inputs at step $j$. If $j=j^{*}$, then $\mathcal{B}$ has found a collision. Therefore,
$\mathbf{Adv}^{\mathrm{CR}}_{H,\mathcal{B}}(\lambda)\;\geq\;\frac{1}{n}\cdot\mathbf{Adv}^{\mathrm{NE}}_{\Pi,\mathcal{A}}(\lambda),$
which gives the claimed bound. ∎
5.3 Discussion
The three theorems above show that our construction reduces the security of constant-size evidence structures to standard, well-studied assumptions: collision resistance of the hash function $H$ and existential unforgeability of the signature scheme under chosen-message attacks. The reductions are tight up to small polynomial factors ($k$ for the field count and $n$ for the chain length), which are concrete and small in practice (typically $k\leq 12$ and $n\leq 10^{6}$).
These results provide rigorous foundations for the security of our construction, reducing all security properties to standard, well-studied cryptographic assumptions. We note that extensions to adaptive corruption models (where the adversary may compromise signing keys during the protocol), concurrent multi-chain sessions, and stronger notions such as forward security [4] remain interesting directions for future work.